### v1

@@ Abstract
We study the finite element approximation of problems involving the weighted $p$-Laplacian for $p\in(1,\infty)$ and weights belonging to the Muckenhoupt class $A_{1}$. In particular, we consider an equation and an obstacle problem for the weighted $p$-Laplacian and derive error estimates in both cases. The analysis is based on the language of weighted Orlicz and Orlicz–Sobolev spaces.
@@
nonlinear elliptic equations, weighted $p$-Laplacian, obstacle problem, Muckenhoupt weights, finite element discretizations, quasi norms, a priori error estimates.
@@ 1 Introduction
Let $d\in\mathbb{N}$ with $d\geq 2$, and let $\Omega$ be a bounded polytope in $\mathbb{R}^{d}$ with Lipschitz boundary $\partial\Omega$. The aim of this paper is to study finite element methods for approximating the unconstrained and constrained minimization of the following convex energy functional
(1) $\mathcal{J}(v)\coloneqq\int_{\Omega}\omega(x)\varphi(|\nabla v(x)|)\mathrm{d}x-\int_{\Omega}\omega(x)f(x)v(x)\mathrm{d}x,$
over an appropriate function class that satisfies homogeneous Dirichlet boundary conditions in particular. Here $\varphi$ is a suitable convex function, $\omega$ is a weight belonging to a Muckenhoupt class, and $f$ is a given forcing term; see Section 2 for precise assumptions.
@@
the vector field $\boldsymbol{\mathcal{A}}$ may have a nonstandard growth (through the function $\varphi$) and also a non-translation invariant property (due to the dependence on the point $x$). The only simplifying assumption we make is that we consider each of these phenomena separately in the sense that
$\boldsymbol{\mathcal{A}}(x,\boldsymbol{\zeta})=\omega(x){\mathbf{A}}(\boldsymbol{\zeta}),\qquad\mathbf{A}(\boldsymbol{\zeta})=\varphi^{\prime}(|\boldsymbol{\zeta}|)\frac{\boldsymbol{\zeta}}{|\boldsymbol{\zeta}|}$
@@
The analysis and numerical approximation of quasilinear problems in general and the $p$-Laplacian in particular have a rich history, and any attempt to give a complete bibliographical account is doomed to failure. We mention here only the classical references [4] and [17]. In particular, in [17], the authors introduce the language of Orlicz functions to perform an error analysis. For the approximation of obstacle problems with the $p$-Laplacian as operator, we may refer, for instance, to [2, 11].
@@
We structure our presentation as follows. Section 2 introduces the notation and gives an overview of the main ingredients we will use. The goal is to properly describe the functional framework we will use: weighted Orlicz and Orlicz–Sobolev spaces. This will be done in Section 2.5. The unconstrained minimization of (1) and its numerical approximation will be discussed in Section 3. The analysis hinges on the properties of the Scott-Zhang operator on weighted Orlicz spaces. Finally, an obstacle problem related to (1) and its numerical approximation is studied in Section 4. In this case, the analysis relies on a positivity preserving interpolant and its properties. This operator is analyzed in Section 4.3.
@@ 2 Notation and preliminary remarks
The first notation we introduce is the relation $A\lesssim B$. This means that $A\leq CB$ for a nonessential constant $C>0$, which can change at each occurrence. The relation $A\gtrsim B$ means $B\lesssim A$, and $A\simeq B$ is the short form for $A\lesssim B\lesssim A$. If it is necessary to explicitly mention a constant $C$, we assume that $C>0$ and that the value can change at each occurrence.
Let $E\subset\mathbb{R}^{d}$ be a measurable set. We denote by $|E|$ the $d$-dimensional Lebesgue measure of the set $E$. If $0<|E|<\infty$ and $f\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{d})$, we define the mean value of the function $f$ over the set $E$ by
$\langle f\rangle_{E}\coloneqq\fint_{E}f(x)\,\mbox{\rm d}x=\frac{1}{|E|}\int_{E}f(x)\,\mbox{\rm d}x.$
@@ 2.1 N-functions, complementary functions, and shifted N-functions
We say that $\Phi:[0,\infty)\rightarrow[0,\infty)$ is an $N$-function if $\Phi$ is differentiable and its derivative $\Phi^{\prime}$ satisfies the following properties: $\Phi^{\prime}(0)=0$, $\Phi^{\prime}(s)>0$ for $s>0$, $\Phi^{\prime}$ is right-continuous at any point $s\geq 0$, $\Phi^{\prime}$ is nondecreasing on $[0,\infty)$, and $\lim_{s\rightarrow\infty}\Phi^{\prime}(s)=\infty$ [39, Definition 4.2.1]. We note that every $N$-function is convex [39, Lemma 4.2.2], [33, page 7]. We say that an $N$-function $\Phi$ satisfies the $\Delta_{2}$-condition (we shall write $\Phi\in\Delta_{2}$) if there exists a positive constant $K$ such that $\Phi(2t)\leq K\Phi(t)$ for all $t\geq 0$. The smallest of these constants is denoted by $\Delta_{2}(\Phi)$ [39, Definition 4.4.1], [33, page 23]. Since $\Phi(t)\leq\Phi(2t)$, the $\Delta_{2}$-condition guarantees that $\Phi(t)\simeq\Phi(2t)$ for all $t\geq 0$. More generally, if $\Delta_{2}(\Phi)<\infty$ and $a>1$ is fixed, then $\Phi(t)\simeq\Phi(at)$ for all $t\geq 0$ [39, Exercise 4.5.5], [33, inequality (4.2)]. Finally, we mention that if $\Phi$ is an $N$-function and $\Delta_{2}(\Phi)<\infty$, then $\Phi(t)\simeq\Phi^{\prime}(t)\,t$ uniformly in $t\geq 0$ [34, Proposition 7(2.6b)].
@@ 2.1.1 Complementary functions
Let $\Phi$ be an $N$-function. We set [39, Definition 4.3.1]
$(\Phi^{\prime})^{-1}:[0,\infty)\rightarrow[0,\infty),\qquad(\Phi^{\prime})^{-1}(t)=\sup\{s:\Phi^{\prime}(s)\leq t\}.$
If $\Phi^{\prime}$ is continuous and strictly increasing in $[0,\infty)$, then $(\Phi^{\prime})^{-1}$ is the inverse function of $\Phi^{\prime}$ and vice-versa [39, Remark 4.3.3]. We now define [39, Definition 4.3.1]
$\Phi^{*}:[0,\infty)\rightarrow[0,\infty),\qquad\Phi^{*}(t)=\int_{0}^{t}(\Phi^{\prime})^{-1}(s)\mathrm{d}s.$
@@
The following examples are important in our context. We let $p\in(1,\infty)$ and $\kappa\geq 0$. We define the $N$-functions $\Phi_{p}:[0,\infty)\rightarrow[0,\infty)$ and $\Phi_{p,\kappa}:[0,\infty)\rightarrow[0,\infty)$ as
(3) $\Phi_{p}(t)=\frac{1}{p}t^{p},\qquad\Phi_{p,\kappa}(t)=\int_{0}^{t}\Phi^{\prime}_{p,\kappa}(s)\mathrm{d}s,\qquad\Phi^{\prime}_{p,\kappa}(t)=(\kappa+t)^{p-2}t.$
@@ 2.1.2 Young’s inequalities
If $(\Phi,\Phi^{*})$ is a pair of complementary $N$-functions, then we have
(4) $st\leq\Phi(s)+\Phi^{*}(t)$
for all $s,t\geq 0$. The equality holds if and only if $t=\Phi^{\prime}(s)$ or $s=(\Phi^{\prime})^{-1}(t)$ [39, Theorem 4.3.4]. This is commonly referred to as Young’s inequality. We also have the following refined versions: If $\Phi,\Phi^{*}\in\Delta_{2}$, then for all $\delta>0$ there exists $C_{\delta}>0$, depending on $\delta$, $\Delta_{2}(\Phi)$, and $\Delta_{2}(\Phi^{*})$, such that [15, Lemma 32]
@@
Note that the inequality (4) is a generalization of the classical Young’s inequality: If $p\in(1,\infty)$ and $a,b\geq 0$, then
$ab\leq\frac{1}{p}a^{p}+\frac{1}{p^{\prime}}b^{p^{\prime}},\qquad\frac{1}{p}+\frac{1}{p^{\prime}}=1.$
Similarly, (5) is a generalization of “Young’s inequality with $\delta$”: If $p\in(1,\infty)$ and $a,b\geq 0$, then for all $\delta>0$ there exists $C_{\delta}>0$ such that
$ab\leq\delta a^{p}+C_{\delta}b^{p^{\prime}}.$
@@ 2.1.3 Shifted $N$-functions
(7) $\Phi_{a}:[0,\infty)\rightarrow[0,\infty),\quad\Phi_{a}(t)=\int_{0}^{t}\Phi_{a}^{\prime}(s)\mathrm{d}s,\quad\Phi_{a}^{\prime}(t)\coloneqq\Phi^{\prime}(a+t)\frac{t}{a+t},\,t\geq 0.$

### v2

@@ Abstract
We study the finite element approximation of problems involving the weighted $\Phi$-Laplacian, where $\Phi$ is an $N$-function and the weight belongs to the class $A_{\Phi}$. In particular, we consider a boundary value problem and an obstacle problem and derive error estimates in both cases. The analysis is based on the language of weighted Orlicz and weighted Orlicz–Sobolev spaces.
@@
nonlinear elliptic equations, weighted $\Phi$-Laplacian, obstacle problem, Muckenhoupt weights, finite element discretizations, quasi norms, a priori error estimates.
@@ 1 Introduction
Let $d\in\mathbb{N}$ with $d\geq 2$, and let $\Omega$ be a bounded polytope in $\mathbb{R}^{d}$ with Lipschitz boundary $\partial\Omega$. We study finite element methods for approximating the unconstrained and constrained minimization of the following convex energy functional
(1) $\mathcal{J}(v)\coloneqq\int_{\Omega}\omega(x)\Phi(|\nabla v(x)|)\mathrm{d}x-\int_{\Omega}\omega(x)f(x)v(x)\mathrm{d}x,$
over an appropriate function class that, in particular, satisfies homogeneous Dirichlet boundary conditions. Here $\Phi$ is a suitable convex function, $\omega$ is a weight belonging to a suitable class adapted to $\Phi$, and $f$ is a given forcing term; see Section 2 for details.
@@
the vector field $\boldsymbol{\mathcal{A}}$ may have a nonstandard growth through $\Phi$ and is not translation invariant due to the dependence on the point $x$. The only simplifying assumption we make is that we consider each of these phenomena separately in the sense that
(2) $\boldsymbol{\mathcal{A}}(x,\boldsymbol{\zeta})=\omega(x){\mathbf{A}}(\boldsymbol{\zeta}),\qquad\mathbf{A}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},\qquad\mathbf{A}(\boldsymbol{\zeta})=\Phi^{\prime}(|\boldsymbol{\zeta}|)\frac{\boldsymbol{\zeta}}{|\boldsymbol{\zeta}|}$
@@
The analysis and numerical approximation of quasilinear problems in general and the $p$-Laplacian in particular have a rich history, and any attempt to give a complete bibliographical account is doomed to failure. We mention here only the classical references [4] and [17]. In particular in [17] the authors introduce the language of Orlicz functions to perform an error analysis. For the approximation of obstacle problems with the $p$-Laplacian as operator, we can refer, for instance, to [2, 11].
@@
Our presentation is as follows. Section 2 introduces notation and gives an overview of the main ingredients we will use. The goal is to properly describe the functional framework we use: weighted Orlicz and Orlicz–Sobolev spaces. The unconstrained minimization of (1) and its numerical approximation are discussed in Sections 3 and 4, respectively. The analysis hinges on the properties of the Scott-Zhang operator on weighted Orlicz spaces. Finally, an obstacle problem related to (1) and its numerical approximation are studied in Section 5. In this case, the analysis relies on a positivity preserving interpolant and its properties. This operator is analyzed in Section 5.3.
@@ 2 Notation and preliminary remarks
The first notation we introduce is the relation $A\lesssim B$. This means that $A\leq CB$ for a nonessential constant $C>0$, which can change at each occurrence. The relation $A\gtrsim B$ means $B\lesssim A$, and $A\simeq B$ is the short form for $A\lesssim B\lesssim A$. If it is necessary to explicitly mention a constant $C$, we assume that $C>0$ and that its value can change at each occurrence.
Let $E\subset\mathbb{R}^{d}$ be a measurable set, and let $|E|$ be the $d$-dimensional Lebesgue measure of $E$. If $0<|E|<\infty$ and $v\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{d})$, we denote the mean value of $v$ over the set $E$ by
$\langle v\rangle_{E}\coloneqq\fint_{E}v(x)\,\mbox{\rm d}x=\frac{1}{|E|}\int_{E}v(x)\,\mbox{\rm d}x.$
@@ 2.1 $N$-functions, complementary functions, and shifted $N$-functions
We say that $\Phi:[0,\infty)\rightarrow[0,\infty)$ is an $N$-function if $\Phi$ is differentiable and its derivative $\Phi^{\prime}$ satisfies the following properties: $\Phi^{\prime}(0)=0$, $\Phi^{\prime}(s)>0$ for $s>0$, $\Phi^{\prime}$ is right-continuous at any point $s\geq 0$, $\Phi^{\prime}$ is nondecreasing on $0,\infty)$, and $\lim_{s\rightarrow\infty}\Phi^{\prime}(s)=\infty$ [42, Definition 4.2.1]. We note that every $N$-function is convex [42, Lemma 4.2.2], [34, page 7]. We say that an $N$-function $\Phi$ satisfies the $\Delta_{2}$-condition (we shall write $\Phi\in\Delta_{2}$) if there exists a positive constant $K$ such that $\Phi(2t)\leq K\Phi(t)$ for all $t\geq 0$. The smallest of these constants is denoted by $\Delta_{2}(\Phi)$ [42, Definition 4.4.1], [34, page 23]. Since $\Phi(t)\leq\Phi(2t)$, the $\Delta_{2}$-condition guarantees that $\Phi(t)\simeq\Phi(2t)$ for all $t\geq 0$. More generally, if $\Delta_{2}(\Phi)<\infty$ and $a>1$ is fixed, then $\Phi(t)\simeq\Phi(at)$ for all $t\geq 0$ [42, Exercise 4.5.5], [34, inequality (4.2)]. Finally, we mention that if $\Phi$ is an $N$-function and $\Delta_{2}(\Phi)<\infty$, then $\Phi(t)\simeq\Phi^{\prime}(t)\,t$ uniformly in $t\geq 0$ [37, Proposition 7(2.6b)].
To characterize the behavior of an $N$-function $\Phi$ for very small and very large values of its argument, we introduce the lower and upper indices, respectively, as [33, Definition 1.1.5], [14, page 71]
(4) $i(\Phi)\coloneqq\lim_{\lambda\downarrow 0}\frac{1}{\log\lambda}\log\left(\sup_{t>0}\frac{\Phi(\lambda t)}{\Phi(t)}\right),\qquad I(\Phi)\coloneqq\lim_{\lambda\uparrow\infty}\frac{1}{\log\lambda}\log\left(\sup_{t>0}\frac{\Phi(\lambda t)}{\Phi(t)}\right).$
It follows from the definition that $1\leq i(\Phi)\leq I(\Phi)\leq\infty$. Moreover, if $\Phi\in\Delta_{2}$, then $1\leq i(\Phi)\leq I(\Phi)<\infty$ [14, page 71]. Finally, we have that for every $\epsilon>0$ there are constants $c_{\epsilon}$ and $C_{\epsilon}$ satisfying $0<c_{\epsilon}\leq C_{\epsilon}$ such that, for every $s,t\geq 0$ [36, page 114],
(5) $c_{\epsilon}\min\{t^{i(\Phi)-\epsilon},t^{I(\Phi)+\epsilon}\}\Phi(s)\leq\Phi(st)\leq C_{\epsilon}\max\{t^{i(\Phi)-\epsilon},t^{I(\Phi)+\epsilon}\}\Phi(s).$
@@ 2.1.1 Complementary functions
Let $\Phi$ be an $N$-function. Define $(\Phi^{\prime})^{-1}:[0,\infty)\rightarrow[0,\infty)$ such that $(\Phi^{\prime})^{-1}(t)\coloneqq\sup\{s:\Phi^{\prime}(s)\leq t\}$ [42, Definition 4.3.1]. If $\Phi^{\prime}$ is continuous and strictly increasing in $0,\infty)$, then $(\Phi^{\prime})^{-1}$ is the inverse function of $\Phi^{\prime}$ and vice-versa [42, Remark 4.3.3]. We now define [42, Definition 4.3.1]
$\Phi^{*}:[0,\infty)\rightarrow[0,\infty),\qquad\Phi^{*}(t)\coloneqq\int_{0}^{t}(\Phi^{\prime})^{-1}(s)\mathrm{d}s.$
@@
Let $p\in(1,\infty)$ and $\kappa\geq 0$. We define the $N$-functions $\Phi_{p}:[0,\infty)\rightarrow[0,\infty)$ and $\Phi_{p,\kappa}:[0,\infty)\rightarrow[0,\infty)$ as
(6) $\Phi_{p}(t)\coloneqq\frac{1}{p}t^{p},\qquad\Phi_{p,\kappa}(t)\coloneqq\int_{0}^{t}\Phi^{\prime}_{p,\kappa}(s)\mathrm{d}s,\qquad\Phi^{\prime}_{p,\kappa}(t)\coloneqq(\kappa+t)^{p-2}t.$
@@ 2.1.2 Young’s inequalities
If $(\Phi,\Phi^{*})$ are complementary $N$-functions, then
(7) $st\leq\Phi(s)+\Phi^{*}(t),\qquad\forall s,t\geq 0.$
The equality holds if and only if $t=\Phi^{\prime}(s)$ or $s=(\Phi^{\prime})^{-1}(t)$ [42, Theorem 4.3.4]. This is commonly referred to as Young’s inequality. We also have the following refined versions: If $\Phi,\Phi^{*}\in\Delta_{2}$, then for all $\delta>0$ there exists $C_{\delta}>0$, depending on $\delta$, $\Delta_{2}(\Phi)$, and $\Delta_{2}(\Phi^{*})$ such that [15, Lemma 32]
@@
Note that inequality (7) is a generalization of the classical Young’s inequality, whereas (8) is a generalization of “Young’s inequality with $\delta$”.
@@ 2.1.3 Shifted $N$-functions
(10) $\Phi_{a}:[0,\infty)\rightarrow[0,\infty),\quad\Phi_{a}(t)\coloneqq\int_{0}^{t}\Phi_{a}^{\prime}(s)\mathrm{d}s,\quad\Phi_{a}^{\prime}(t)\coloneqq\Phi^{\prime}(a+t)\frac{t}{a+t},\,t\geq 0.$